# Data-Free Hard-Label Robustness Stealing Attack

### 10 Dec 2023 v1

@@ Abstract
The popularity of Machine Learning as a Service (MLaaS) has led to increased concerns about Model Stealing Attacks (MSA), which aim to craft a clone model by querying MLaaS. Currently, most research on MSA assumes that MLaaS can provide soft labels and that the attacker has a proxy dataset with a similar distribution. However, this fails to encapsulate the more practical scenario where only hard labels are returned by MLaaS and the data distribution remains elusive. Furthermore, most existing work focuses solely on stealing the model accuracy, neglecting the model robustness, while robustness is essential in security-sensitive scenarios, e.g., face-scan payment. Notably, improving model robustness often necessitates the use of expensive techniques such as adversarial training, thereby further making stealing robustness a more lucrative prospect. In response to these identified gaps, we introduce a novel Data-Free Hard-Label Robustness Stealing (DFHL-RS) attack in this paper, which enables the stealing of both model accuracy and robustness by simply querying hard labels of the target model without the help of any natural data. Comprehensive experiments demonstrate the effectiveness of our method. The clone model achieves a clean accuracy of 77.86% and a robust accuracy of 39.51% against AutoAttack, which are only 4.71% and 8.40% lower than the target model on the CIFAR-10 dataset, significantly exceeding the baselines. Our code is available at: https://github.com/LetheSec/DFHL-RS-Attack
Introduction
@@
Furthermore, numerous security-sensitive scenarios require deployed models are not only accurate but also robust to various attacks, such as adversarial attacks. To address this issue, MLaaS providers can employ adversarial training (AT) techniques (Madry et al. 2017) to improve the robustness of their models (Goodman and Xin 2020; Shafique et al. 2020). Despite the target model’s robustness, most existing MSA are limited to Accuracy Stealing, i.e., reconstructing a model with similar accuracy to the target model, and fail at Robustness Stealing, i.e., acquiring the adversarial robustness of the target model while maintaining accuracy. Since the improvement of model robustness requires much more computational resources and extra data (Schmidt et al. 2018; Gowal et al. 2021), robustness stealing will bring greater losses to MLaaS providers. Moreover, if an attacker seeks to train a clone model for transfer-based adversarial attacks against a robust target model, then it becomes crucial to employ robustness stealing to achieve effective attack performance (Dong et al. 2020; Gao et al. 2020).
@@
where $\mathcal{L}_{CE}$ is the cross-entropy loss function, $\theta_{M_{C}}$ represents parameters of the clone model, and $x_{q}+\delta$ represents the adversarial examples generated using PGD (Madry et al. 2017). However, this method will also inherit the shortcomings of AT, which will seriously reduce the clean accuracy.
Uncertain Examples (UE)
@@
High-Entropy Examples (HEE)
@@
Comparison of Different Query Samples
@@
Data-Free Hard-Label Robustness Stealing
@@
Substitute Data Generation
@@
Clone Model Training
@@
We consider two benchmark datasets commonly used in AT research (Madry et al. 2017; Zhang et al. 2019; Li et al. 2023b), CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al. 2009), as target datasets. Prior work requires a proxy dataset with samples from the same distribution (Tramèr et al. 2016; Jagielski et al. 2020; Orekondy, Schiele, and Fritz 2019) or the same task domain (Sanyal, Addepalli, and Babu 2022; Li et al. 2023a) as the target dataset. However, we avoid using any natural data.
@@
We evaluate our attack on different models with various architectures. $M_{T}$ is selected from ResNet18 (He et al. 2016) and WideResNet-34-10 (Zagoruyko and Komodakis 2016). $M_{C}$ may be different from $M_{T}$ in architecture, so we use two additional models, i.e., ResNet34 and MobileNetV2 (Sandler et al. 2018). We employ two commonly used AT strategies, namely PGD-AT (Madry et al. 2017) and TRADES (Zhang et al. 2019), and a state-of-the-art method, STAT-AWP (Li et al. 2023b), to improve the robustness of $M_{T}$. We follow the same generator architecture as (Fang et al. 2021, 2022).
@@
In this task, $M_{C}$ should not only have good usability, but also need to have certain robustness. Therefore, in addition to considering clean accuracy, i.e., the accuracy over clean samples, we also measure robust accuracy against various adversarial attacks, including FGSM (Goodfellow, Shlens, and Szegedy 2014), PGD-20, PGD-100 (Madry et al. 2017), CW-100 (Carlini and Wagner 2017) and AutoAttack (AA) (Croce and Hein 2020). The attack settings are $\epsilon=8/255$ and $\eta=2/255$. The number of attack step is 20 for PGD-20, and 100 for PGD-100 and CW-100.
@@ Acknowledgements
This work was supported in part by the Natural Science Foundation of China under Grant U20B2047, 62102386, U2336206, 62072421, 62002334 and 62121002, and by Xiaomi Young Scholars Program.
Appendix
Details of Figure 1
@@ 20             Update $\theta_{C}\leftarrow\eta_{C}\nabla_{\theta_{C}}\mathcal{L}_{C}$
Performance of Target Models
@@
More Ablation Results of Query Budget
@@
Taxonomy of Prior Work
@@
Details of Data Augmentation

### 12 Dec 2023 v2

@@ Abstract
The popularity of Machine Learning as a Service (MLaaS) has led to increased concerns about Model Stealing Attacks (MSA), which aim to craft a clone model by querying MLaaS. Currently, most research on MSA assumes that MLaaS can provide soft labels and that the attacker has a proxy dataset with a similar distribution. However, this fails to encapsulate the more practical scenario where only hard labels are returned by MLaaS and the data distribution remains elusive. Furthermore, most existing work focuses solely on stealing the model accuracy, neglecting the model robustness, while robustness is essential in security-sensitive scenarios, e.g., face-scan payment. Notably, improving model robustness often necessitates the use of expensive techniques such as adversarial training, thereby further making stealing robustness a more lucrative prospect. In response to these identified gaps, we introduce a novel Data-Free Hard-Label Robustness Stealing (DFHL-RS) attack in this paper, which enables the stealing of both model accuracy and robustness by simply querying hard labels of the target model without the help of any natural data. Comprehensive experiments demonstrate the effectiveness of our method. The clone model achieves a clean accuracy of 77.86% and a robust accuracy of 39.51% against AutoAttack, which are only 4.71% and 8.40% lower than the target model on the CIFAR-10 dataset, significantly exceeding the baselines. Our code is available at: https://github.com/LetheSec/DFHL-RS-Attack.
1 Introduction
@@
Furthermore, numerous security-sensitive scenarios require deployed models are not only accurate but also robust to various attacks, such as adversarial attacks. To address this issue, MLaaS providers can employ adversarial training (AT) techniques (Madry et al. 2018) to improve the robustness of their models (Goodman and Xin 2020; Shafique et al. 2020). Despite the target model’s robustness, most existing MSA are limited to Accuracy Stealing, i.e., reconstructing a model with similar accuracy to the target model, and fail at Robustness Stealing, i.e., acquiring the adversarial robustness of the target model while maintaining accuracy. Since the improvement of model robustness requires much more computational resources and extra data (Schmidt et al. 2018; Gowal et al. 2021), robustness stealing will bring greater losses to MLaaS providers. Moreover, if an attacker seeks to train a clone model for transfer-based adversarial attacks against a robust target model, then it becomes crucial to employ robustness stealing to achieve effective attack performance (Dong et al. 2020; Gao et al. 2020).
@@ 3.1 Adversarial Training (AT)
where $\mathcal{L}_{CE}$ is the cross-entropy loss function, $\theta_{M_{C}}$ represents parameters of the clone model, and $x_{q}+\delta$ represents the adversarial examples generated using PGD (Madry et al. 2018). However, this method will also inherit the shortcomings of AT, which will seriously reduce the clean accuracy.
3.2 Uncertain Examples (UE)
@@
3.3 High-Entropy Examples (HEE)
@@ 3.4 Adversarial Examples (AE)
3.5 Comparison of Different Query Samples
@@
4 Data-Free Hard-Label Robustness Stealing
@@ 4.1 Overview
4.2 Substitute Data Generation
@@
4.3 Clone Model Training
@@ 5.1 Experimental Settings
We consider two benchmark datasets commonly used in AT research (Madry et al. 2018; Zhang et al. 2019; Li et al. 2023b), CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al. 2009), as target datasets. Prior work requires a proxy dataset with samples from the same distribution (Tramèr et al. 2016; Jagielski et al. 2020; Orekondy, Schiele, and Fritz 2019) or the same task domain (Sanyal, Addepalli, and Babu 2022; Li et al. 2023a) as the target dataset. However, we avoid using any natural data.
@@
We evaluate our attack on different models with various architectures. $M_{T}$ is selected from ResNet18 (He et al. 2016) and WideResNet-34-10 (Zagoruyko and Komodakis 2016). $M_{C}$ may be different from $M_{T}$ in architecture, so we use two additional models, i.e., ResNet34 and MobileNetV2 (Sandler et al. 2018). We employ two commonly used AT strategies, namely PGD-AT (Madry et al. 2018) and TRADES (Zhang et al. 2019), and a state-of-the-art method, STAT-AWP (Li et al. 2023b), to improve the robustness of $M_{T}$. We follow the same generator architecture as (Fang et al. 2021, 2022).
@@
In this task, $M_{C}$ should not only have good usability, but also need to have certain robustness. Therefore, in addition to considering clean accuracy, i.e., the accuracy over clean samples, we also measure robust accuracy against various adversarial attacks, including FGSM (Goodfellow, Shlens, and Szegedy 2014), PGD-20, PGD-100 (Madry et al. 2018), CW-100 (Carlini and Wagner 2017) and AutoAttack (AA) (Croce and Hein 2020). The attack settings are $\epsilon=8/255$ and $\eta=2/255$. The number of attack step is 20 for PGD-20, and 100 for PGD-100 and CW-100.
@@ Acknowledgements
This work was supported in part by the Natural Science Foundation of China under Grant U20B2047, 62102386, U2336206, 62072421 and 62121002, and by Xiaomi Young Scholars Program.
Appendix A Appendix
A.1 Details of Figure 1
@@ A.3 The Impact of Different $N_{G}$
A.4 Performance of Target Models
@@
A.5 More Ablation Results of Query Budget
@@
A.6 Taxonomy of Prior Work
@@
A.7 Details of Data Augmentation